# Strict Optimality of Frequency Estimation Under Local Differential Privacy

### 12 Mar 2026 v1

@@ 7 Conclusion
This paper establishes the strict optimality of frequency estimation under LDP by proving that the optimal symmetric configuration achieves the strict lower bound for both $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ losses. We further demonstrate that the communication cost of an optimal frequency estimator can be reduced to $\log_{2}\left(\frac{d(d-1)}{2}+1\right)$, where $d$ is the dictionary size, and we develop an algorithm to construct an optimal estimator within this communication bound. Consequently, our results offer a practical deployment strategy: Optimized Count-Mean Sketch is recommended for large dictionaries, while weighted and original Subset Selection are ideal for smaller scales. Experimental evaluations confirm that these optimal algorithms align perfectly with our derived strict lower bounds.
Appendix A Proof regarding Symmetric Configuration
@@
A.2 Proof of Theorem 2.3 and Corollary 2.2
Reusing the notation from Appendix A.1, the output frequency is expressed as $\mathbf{f_{O}}=\frac{1}{n}\sum_{x\in[d]}Y_{x}=\frac{1}{n}\sum_{x\in[d]}\text{MN}(nf(x),P_{*,x})$, where MN denotes a multinomial distribution and $P$ is the perturbation matrix. Given that $\hat{f}(x)=Q_{x,*}\mathbf{f_{O}}$, where $Q_{x,*}$ is the $x$-th row of $Q$, we have:
@@ Appendix B Proof of Precision Optimality
B.3 The proof of Lemma 3.1
Let $d$, $\epsilon$, and $n$ denote the dictionary size, the privacy factor, and the dataset size, respectively. Suppose we have a LDP frequency estimator that satisfies the optimal symmetric configuration with support size $k$. Let there be $m$ possible output responses from its perturbation, denoted by $o\in[m]$. We define $p_{o}$ as the base probability of output $o$, and $S_{o,x}\in\{0,1\}$ as the indicator that $o$ supports $x\in[d]$. For a symmetric configuration, the parameters $p^{*}$ and $q^{*}$ can be expressed as:
@@
B.4 Proof of Theorem 3.4
[…]
$\mathcal{L}_{2}(\hat{f})=\max_{\forall\mathbf{X}:|\mathbf{X}|=n}d\frac{k(e^{\epsilon}-1)+d-1][k(e^{\epsilon}-1)+d-e^{\epsilon}]}{nk(d-k)(e^{\epsilon}-1)^{2}}\\ +\frac{k^{2}(e^{\epsilon}-1)+k(2d-e^{\epsilon}-1)+d-d^{2}}{nk(d-k)(e^{\epsilon}-1)}$ (21)
Since the parameters are independent of any $\mathbf{X}$ with size $n$, we can remove the maximum operator. Differentiating Eq. (21) with respect to $k$, we acquire
$\frac{\partial\mathcal{L}_{2}(\hat{f})}{\partial k}=\frac{(d-1)^{2}(ke^{\epsilon}-k+d)(ke^{\epsilon}+k-d)}{k^{2}(k-d)^{2}(e^{\epsilon}-1)^{2}}.$
@@
Recall that an optimized frequency estimator requires an optimal symmetric configuration with an optimized support size (Definition 3.1).
@@ Appendix D Proof of Optimized Count Mean Sketch
The operation $h^{\prime}(x)\mod{m}$ then maps these permuted elements into $m$ buckets. Since $d^{\prime}$ is not divisible by $m$, some responses support $\lceil\frac{d^{\prime}}{m}\rceil$ elements while others support $\lfloor\frac{d^{\prime}}{m}\rfloor$ elements.
@@
The support matrix $S$ of OCMS consists of $d^{\prime}(d^{\prime}-1)m$ rows. Given that the hash function $h$ is sampled uniformly and the randomized response has a constant base probability, the base probability of SS is also constant. Although Lemma B.1 assumes a constant support size, it generalizes to any symmetric configuration where $k=\frac{\sum_{o}p_{o}k_{o}}{\sum_{o}p_{o}}$. For OCMS, since $k=d^{\prime}/m$, the base probability is $p_{o}=\frac{d^{\prime}}{\frac{d^{\prime}}{m}(e^{\epsilon}-1)+d^{\prime}}\cdot\frac{1}{md^{\prime}(d^{\prime}-1)}$.
We construct the support matrix of the first configuration, $S^{(1)}$, by stacking all rows of $S$ where the response $(h,z)$ satisfies $z\leq r=d^{\prime}\mod m$. Since the buckets corresponding to $z\leq r$ contain $\lceil\frac{d^{\prime}}{m}\rceil$ elements, all rows of $S^{(1)}$ have a constant support size of $\lceil d^{\prime}/m\rceil$. Defining the base probability of $S^{(1)}$ as a constant, Lemma B.1 yields $p_{1}=\frac{d^{\prime}}{(\frac{d^{\prime}-r}{m}+1)(e^{\epsilon}-1)+d^{\prime}}\cdot\frac{1}{rd^{\prime}(d^{\prime}-1)}$.
Similarly, the support matrix of the second configuration, $S^{(2)}$, is constructed from rows where $z>r$, resulting in a constant support size of $\lfloor\frac{d^{\prime}}{m}\rfloor$. Its base probability is computed as $p_{2}=\frac{d^{\prime}}{\frac{d^{\prime}-r}{m}(e^{\epsilon}-1)+d^{\prime}}\cdot\frac{1}{(m-r)d^{\prime}(d^{\prime}-1)}$. It can be verified that $p_{\alpha}p_{1}=p_{0}$ and $(1-p_{\alpha})p_{2}=p_{0}$, demonstrating that the sampling of these two configurations generates the same support matrix $S$.
These split configurations possess constant support sizes and extremal configurations. We now prove that their support matrix and base probability satisfy the requirements of a symmetric configuration. Taking $S^{(1)}$ as an example, this requires $p_{1}\sum_{o}S^{(1)}_{o,i}S^{(1)}_{o,j}$ to be constant for all $i\neq j$. For an arbitrary pair $(i,j)$ where $i\neq j$, the term $\sum_{o}S^{(1)}_{o,i}S^{(1)}_{o,j}$ represents the number of hash functions mapping $(i,j)$ into the same bucket of size $\lceil\frac{d^{\prime}}{m}\rceil$. Elements within such a bucket can form $(\lceil\frac{d^{\prime}}{m}\rceil+1)(\lceil\frac{d^{\prime}}{m}\rceil)$ distinct pairs, and there are $r$ such buckets. Mapping to the same bucket is equivalent to mapping to these distinct pairs. By Lemma D.1, there are exactly $r(\lceil\frac{d^{\prime}}{m}\rceil+1)(\lceil\frac{d^{\prime}}{m}\rceil)$ hash functions mapping $(i,j)$ to these pairs. This holds for every $(i,j)$, ensuring the sum is constant.
@@
Furthermore, extending this counting method to $m$ buckets and ignoring the remainder $r$, one can derive $P(h(x_{1})=h(x_{2}))\approx\frac{d^{\prime}-m}{m(d^{\prime}-1)}$, verifying the reduction in collision probability discussed in Section 5.2.
Finally, if a process generates variable $A$ with probability $p$ and variable $B$ with probability 1p1-p, and $E[A]=E[B]$, the variance of the process is $p\text{Var}(A)+(1-p)\text{Var}(B)$. This proves Corollary 5.1.
@@
where $\text{Var}(\hat{f}(x)\mid d^{\prime})$ denotes the variance calculated with $d^{\prime}$ in place of $d$. Assuming an optimized frequency estimator (with the $k$ term addressed in the subsequent step), we set $k=\frac{d^{\prime}}{e^{\epsilon}+1}$ to optimize $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. Given Eq. (21), we obtain:
@@
Next, we evaluate the deviation of $k$ from the optimal value $k^{*}$ for a given dictionary size $d^{\prime}$. If $m=\frac{d^{\prime}}{k^{*}}$ is an integer, OCMS is equivalent to sampling between the configurations with support size of $\lceil k^{*}\rceil$ and $\lfloor k^{*}\rfloor$. If $m=\frac{d^{\prime}}{k^{*}}$ is not an integer, then $m$ is either $\lceil\frac{d^{\prime}}{k^{*}}\rceil$ or $\lfloor\frac{d^{\prime}}{k^{*}}\rfloor$, which correspond to sampling between configurations with support size of either ($k^{*}-1$ and $k^{*}$) or ($k^{*}+1$ and $k^{*}$). In all such scenarios, the support size of the sampled configurations deviates from $k^{*}$ by at most one. Analogous to Eq. (22), we define the relative change in variance as:

### 22 Mar 2026 v2

@@ 7 Conclusion
This paper establishes the strict optimality of frequency and distribution estimation under LDP by proving that the optimal symmetric configuration achieves the theoretical lower bound of $\mathcal{L}_{2}$ loss. The theoretical $\mathcal{L}_{1}$ lower bound is also achieved asymptotically. We further demonstrate that the communication cost of an optimal estimator can be reduced to $\log_{2}\left(\frac{d(d-1)}{2}+1\right)$ bits, where $d$ is the dictionary size, and we develop an algorithm to construct an optimal estimator within this communication bound. Consequently, our results offer a practical deployment strategy: Optimized Count-Mean Sketch is recommended for large dictionaries, while weighted and original Subset Selection are ideal for smaller scales. Experimental evaluations confirm that these optimal algorithms align perfectly with our derived strict lower bounds.
Appendix A Proof regarding Linear Estimator
@@
A.2 Proof of Theorem 2.4
Every $M(X^{(i)})$ is sampled i.i.d. from a category distribution with probability $P\bm{\theta}$, which derives that
$\text{Cov}(\sum M(X^{(i)}))=n[\text{diag}(P\bm{\theta})-P\bm{\theta}\bm{\theta}^{T}P^{T}].$
Following Appendix A.1, we prove that
$\text{Var}(Q\mathbf{f_{O}})=(Q\circ Q)P\bm{\theta}-\bm{\theta}^{2}.\qed$
A.3 Proof of Theorem 2.3 and Corollary 2.1
This proof focuses on frequency estimation, but it is applicable to distribution estimation. Reusing the notation from Appendix A.1, the output frequency is expressed as $\mathbf{f_{O}}=\frac{1}{n}\sum_{x\in[d]}Y_{x}=\frac{1}{n}\sum_{x\in[d]}\text{MN}(nf(x),P_{*,x})$, where MN denotes a multinomial distribution and $P$ is the perturbation matrix. Given that $\hat{f}(x)=Q_{x,*}\mathbf{f_{O}}$, where $Q_{x,*}$ is the $x$-th row of $Q$, we have:
@@ Appendix B Proof of Precision Optimality
Let $\bm{\theta}\in\mathbb{R}^{d}$ denote the parameter vector of a categorical distribution $\mathcal{C}(\bm{\theta})$. Define the random variable $O^{(i)}$ as the $i$-th output of $M(\mathcal{C}(\bm{\theta}))$, where mechanism $M$ has a perturbation matrix $P$. The probability of observing output value $o$ is given by $\Pr(O^{(i)}=o)=\mathbf{e}_{o}^{T}P\bm{\theta}$, where $\mathbf{e}_{o}$ is the $o$-th unit basis vector. The score function $D(O^{(i)})$ is defined as:
$D(O^{(i)})$ $=\nabla_{\bm{\theta}}\log(\Pr(O^{(i)}))=\frac{\nabla_{\bm{\theta}}(\mathbf{e}_{O^{(i)}}^{T}P\bm{\theta})}{\Pr(O^{(i)})}$
$=\frac{P^{T}\mathbf{e}_{O^{(i)}}}{\Pr(O^{(i)})}=P^{T}\text{diag}(P\bm{\theta})^{-1}\mathbf{e}_{O^{(i)}}.$
The expectation of the score function is:
$\mathbb{E}[D(O^{(i)})]=\sum_{o\in[m]}\Pr(o)\frac{P^{T}\mathbf{e}_{o}}{\Pr(o)}\\ =P^{T}\sum_{o\in[m]}\mathbf{e}_{o}=P^{T}\mathbf{1}_{m}=\mathbf{1}_{d}.$ (22)
The second moment of the unit vectors is given by:
$\mathbb{E}[\mathbf{e}_{O^{(i)}}\mathbf{e}_{O^{(i)}}^{T}]=\sum_{o\in[m]}\Pr(o)\mathbf{e}_{o}\mathbf{e}_{o}^{T}=\text{diag}(P\bm{\theta}),$
which yields the second moment of the score function:
$\mathbb{E}[D(O^{(i)})D(O^{(i)})^{T}]\\ =P^{T}\text{diag}(P\bm{\theta})^{-1}\mathbb{E}[\mathbf{e}_{O^{(i)}}\mathbf{e}_{O^{(i)}}^{T}]\text{diag}(P\bm{\theta})^{-1}P\\ =P^{T}\text{diag}(P\bm{\theta})^{-1}P.$ (23)
Since $O^{(i)}$ and $O^{(j)}$ are independent for $i\neq j$, their centered score functions are uncorrelated:
$\mathbb{E}[(D(O^{(i)})-\mathbf{1}_{d})(D(O^{(j)})-\mathbf{1}_{d})^{T}]=0.$ (24)
Define the aggregate centered score function as $\overline{D}(\mathbf{O})=\sum_{i=1}^{n}[D(O^{(i)})-\mathbf{1}_{d}]$. From Eqs. (22), (23), and (24), it follows that $\mathbb{E}[\overline{D}(\mathbf{O})]=\mathbf{0}$ and its covariance matrix is:
$\mathbb{E}[\overline{D}(\mathbf{O})\overline{D}(\mathbf{O})^{T}]=n(P^{T}\text{diag}(P\bm{\theta})^{-1}P-\mathbf{1}_{d}\mathbf{1}_{d}^{T})\\ \overset{\text{def}}{=}n(\mathscr{I}_{U}-\mathbf{1}_{d}\mathbf{1}_{d}^{T})\overset{\text{def}}{=}n\mathscr{I}_{C},$ (25)
where $\mathscr{I}_{U}$ and $\mathscr{I}_{C}$ refer to the unconstrained and constrained Fisher Information Matrices (FIM), respectively.
We define $\mathscr{I}_{C}^{+}=\mathscr{I}_{U}^{-1}-\bm{\theta}\bm{\theta}^{T}$ as the generalized inverse of the constrained FIM. Note that:
$\mathscr{I}_{U}\bm{\theta}=P^{T}\text{diag}(P\bm{\theta})^{-1}P\bm{\theta}=P^{T}\mathbf{1}_{m}=\mathbf{1}_{d},$
$\mathscr{I}_{U}^{-1}\mathbf{1}_{d}=\bm{\theta}.$ (26)
Using them with $\bm{\theta}^{T}\mathbf{1}_{d}=1$, we derive:
$\mathscr{I}_{C}^{+}\mathscr{I}_{C}=\mathscr{I}_{U}^{-1}\mathscr{I}_{U}-\mathscr{I}_{U}^{-1}\mathbf{1}_{d}\mathbf{1}_{d}^{T}-\bm{\theta}\bm{\theta}^{T}\mathscr{I}_{U}+\bm{\theta}\bm{\theta}^{T}\mathbf{1}_{d}\mathbf{1}_{d}^{T}\\ =I_{d}-\bm{\theta}\mathbf{1}_{d}-\bm{\theta}\mathbf{1}_{d}+\bm{\theta}\mathbf{1}_{d}^{T}=I_{d}-\bm{\theta}\mathbf{1}_{d},$
which leads to:
$\mathscr{I}_{C}^{+}\mathscr{I}_{C}\mathscr{I}_{C}^{+}=\mathscr{I}_{U}^{-1}-\bm{\theta}\bm{\theta}^{T}-\bm{\theta}\mathbf{1}_{d}^{T}\mathscr{I}_{U}^{-1}+\bm{\theta}\mathbf{1}_{d}\bm{\theta}\bm{\theta}^{T}\\ =\mathscr{I}_{U}^{-1}-\bm{\theta}\bm{\theta}^{T}-\bm{\theta}\bm{\theta}^{T}+\bm{\theta}\bm{\theta}^{T}=\mathscr{I}_{C}^{+}.$ (27)
Let $\hat{\bm{\theta}}$ be the estimated distribution of $\bm{\theta}$ from an unbiased estimator $\hat{\theta}$. We define the cross-covariance matrix $C$ as:
$C$ $=\mathbb{E}[(\hat{\bm{\theta}}-\bm{\theta})\overline{D}(\mathbf{O})^{T}]=\mathbb{E}[\hat{\bm{\theta}}\overline{D}(\mathbf{O})^{T}]$ (28)
$=\mathbb{E}\left[\hat{\bm{\theta}}\sum_{i\in[n]}(\nabla_{\bm{\theta}}\log\Pr(O^{(i)})^{T}-\mathbf{1}^{T})\right]$
$=\sum_{\mathbf{o}\in[m]^{n}}\Pr(\mathbf{o})\hat{\bm{\theta}}\frac{\nabla_{\bm{\theta}}\Pr(\mathbf{o})}{\Pr(\mathbf{o})}-n\mathbb{E}[\hat{\bm{\theta}}\mathbf{1}^{T}]$
$=\nabla_{\bm{\theta}}\sum_{\mathbf{o}\in[m]^{n}}\hat{\bm{\theta}}\Pr(\mathbf{o})-n\bm{\theta}\mathbf{1}^{T}=\nabla_{\bm{\theta}}\mathbb{E}[\hat{\bm{\theta}}]-n\bm{\theta}\mathbf{1}^{T}=I_{d}-n\bm{\theta}\mathbf{1}^{T}.$
Using Eqs. (26) and $\bm{\theta}^{T}\mathbf{1}_{d}=1$, we have:
$C\mathscr{I}_{C}^{+}=\mathscr{I}_{U}^{-1}-n\bm{\theta}\mathbf{1}^{T}\mathscr{I}_{U}^{-1}-\bm{\theta}\bm{\theta}^{T}+n\bm{\theta}\mathbf{1}^{T}\bm{\theta}\bm{\theta}^{T}\\ =\mathscr{I}_{U}^{-1}-n\bm{\theta}\bm{\theta}^{T}-\bm{\theta}\bm{\theta}^{T}+n\bm{\theta}\bm{\theta}^{T}=\mathscr{I}_{C}^{+}.$ (29)
Let $V=\mathbb{E}[(\hat{\bm{\theta}}-\bm{\theta})(\hat{\bm{\theta}}-\bm{\theta})^{T}]$ denote the covariance matrix of $\hat{\bm{\theta}}$. We consider an auxiliary vector $\mathbf{g}=(\hat{\bm{\theta}}-\bm{\theta})-\frac{1}{n}\mathscr{I}_{C}^{+}\overline{D}(\mathbf{O})$. Since the expectation of the outer product of any random vector is positive semi-definite, and given Eqs. (25) and (28), it follows that:
$\mathbb{E}[\mathbf{g}\mathbf{g}^{T}]=V-\frac{1}{n}C\mathscr{I}_{C}^{+}-\frac{1}{n}\mathscr{I}_{C}^{+}C^{T}+\frac{1}{n}\mathscr{I}_{C}^{+}\mathscr{I}_{C}\mathscr{I}_{C}^{+}\succeq 0.$
Using $C\mathscr{I}_{C}^{+}=\mathscr{I}_{C}^{+}$ and $\mathscr{I}_{C}^{+}\mathscr{I}_{C}\mathscr{I}_{C}^{+}=\mathscr{I}_{C}^{+}$ from Eqs. (27) and (29), this simplifies to $V\succeq\frac{1}{n}\mathscr{I}_{C}^{+}$. Therefore:
$L_{2}(\hat{\theta},\bm{\theta},n)$ $=\text{Tr}(V)\geq\frac{1}{n}\text{Tr}(\mathscr{I}_{C}^{+})$
$=\frac{1}{n}\text{Tr}\left[(P^{T}\text{diag}(P\bm{\theta})^{-1}P)^{-1}-\bm{\theta}\bm{\theta}^{T}\right].\qed$
For the $L_{1}$ lower bound, when the dataset size $n$ is asymptotically large, the Hájek-Le Cam Asymptotic Minimax Theorem states that the bowl-shaped loss, e.g., $L_{1}$, of any estimator is lower bounded by that of a normal distribution with covariance $\frac{1}{n}\mathscr{I}_{C}^{+}$, the inverse FIM. For a normal distribution, $\mathbb{E}[|\hat{\theta}_{x}-\theta_{x}|]=\sqrt{\frac{2}{n\pi}\text{Var}(\hat{\theta}_{x})}$. Thus, the $L_{1}$ loss is bounded by:
$\lim_{n\rightarrow\infty}L_{1}(\hat{\theta},\bm{\theta},n)\geq\sum_{x\in[d]}\sqrt{\frac{2}{n\pi}(\mathscr{I}_{C}^{+})_{x,x}}.\qed$
B.2 Proof of Theorem 3.4
For any $x\in[d]$, denote $\hat{\theta}=\hat{f}$ as the $x$-th estimated value of $\hat{\bm{\theta}}=\hat{\mathbf{f}}$, and $f$ and $\theta$ are the $x$-th elements of the sampled frequency $\mathbf{f}$ and the distribution $\bm{\theta}$. We have
$\text{Var}(\hat{\theta})=\mathbb{E}[(\hat{f}-f+f-\theta)^{2}]\\ =\mathbb{E}[(\hat{f}-f)^{2}]+\mathbb{E}[(f-\theta)^{2}]+2\mathbb{E}[(\hat{f}-f)(f-\theta)].$ (30)
Subsequently, we derive
$\mathbb{E}[(\hat{f}-f)^{2}]=\sum_{\mathbf{f}}\Pr(\mathbf{f}|\bm{\theta})\sum_{\hat{f}}\Pr(\hat{f}|\mathbf{f})(\hat{f}-f)^{2}\\ =\sum_{\mathbf{f}}\Pr(\mathbf{f}|\bm{\theta})\text{Var}(\hat{f}|\mathbf{f}).$
Given the category distribution, $\mathbb{E}[(f-\theta)^{2}]=\frac{\theta-\theta^{2}}{n}$. Additionally, given that $\mathbb{E}[\hat{f}-f]=0$
$\mathbb{E}[(\hat{f}-f)(f-\theta)]=\sum_{\mathbf{f}}\Pr(\mathbf{f}|\bm{\theta})\sum_{\hat{f}}\Pr(\hat{f}|\mathbf{f})(\hat{f}-f)(f-\theta)\\ =\sum_{\mathbf{f}}\Pr(\mathbf{f}|\bm{\theta})\mathbb{E}[\hat{f}-f|f](f-\theta)=0.$
Substituting them back to Eq. (30), we obtain
$\text{Var}(\hat{\theta})=\sum_{\mathbf{f}}\Pr(\mathbf{f}|\bm{\theta})\text{Var}(\hat{f}|\mathbf{f})+\frac{\theta-\theta^{2}}{n}.\qed$
B.3 Proof of Theorem 3.6
@@
B.5 The proof of Lemma 3.1
Let $d$, $\epsilon$, and $n$ denote the dictionary size, the privacy parameter, and the dataset size, respectively. Suppose we have a LDP estimator that satisfies the optimal symmetric configuration with support size $k$. Let there be $m$ possible output responses from its perturbation, denoted by $o\in[m]$. We define $p_{o}$ as the base probability of output $o$, and $S_{o,x}\in\{0,1\}$ as the indicator that $o$ supports $x\in[d]$. For a symmetric configuration, the parameters $p^{*}$ and $q^{*}$ can be expressed as:
@@
B.6 Proof of Theorem 3.8
Substituting Eq. (17) into $\mathcal{L}_{2}(\hat{f},n)$, we obtain
$\mathcal{L}_{2}(\hat{f},n)=\max_{\mathbf{f}}d\frac{k(e^{\epsilon}-1)+d-1][k(e^{\epsilon}-1)+d-e^{\epsilon}]}{nk(d-k)(e^{\epsilon}-1)^{2}}\\ +\frac{k^{2}(e^{\epsilon}-1)+k(2d-e^{\epsilon}-1)+d-d^{2}}{nk(d-k)(e^{\epsilon}-1)}$ (36)
Since the parameters are independent of any $\mathbf{f}$, we can remove the maximum operator. Differentiating Eq. (36) with respect to $k$, we acquire
$\frac{\partial\mathcal{L}_{2}(\hat{f},n)}{\partial k}=\frac{(d-1)^{2}(ke^{\epsilon}-k+d)(ke^{\epsilon}+k-d)}{nk^{2}(k-d)^{2}(e^{\epsilon}-1)^{2}}.$
@@
Recall that an optimal estimator requires an optimal symmetric configuration with an optimized support size (Definition 3.2).
@@ Appendix D Proof of Optimized Count Mean Sketch
The operation $h^{\prime}(x)\mod{B}$ then maps these permuted elements into $B$ buckets. Since $d^{\prime}$ is not divisible by $B$, some responses support $\lceil\frac{d^{\prime}}{B}\rceil$ elements while others support $\lfloor\frac{d^{\prime}}{B}\rfloor$ elements.
@@
The support matrix $S$ of OCMS consists of $d^{\prime}(d^{\prime}-1)B$ rows. Given that the hash function $h$ is sampled uniformly and the randomized response has a constant base probability, the base probability of SS is also constant. Although Lemma B.1 assumes a constant support size, it generalizes to any symmetric configuration where $k=\frac{\sum_{o}p_{o}k_{o}}{\sum_{o}p_{o}}$. For OCMS, since $k=d^{\prime}/B$, the base probability is $p_{o}=\frac{d^{\prime}}{\frac{d^{\prime}}{B}(e^{\epsilon}-1)+d^{\prime}}\cdot\frac{1}{Bd^{\prime}(d^{\prime}-1)}$.
We construct the support matrix of the first configuration, $S^{(1)}$, by stacking all rows of $S$ where the response $(h,z)$ satisfies $z\leq r=d^{\prime}\mod B$. Since the buckets corresponding to $z\leq r$ contain $\lceil\frac{d^{\prime}}{B}\rceil$ elements, all rows of $S^{(1)}$ have a constant support size of $\lceil d^{\prime}/B\rceil$. Defining the base probability of $S^{(1)}$ as a constant, Lemma B.1 yields $p_{1}=\frac{d^{\prime}}{(\frac{d^{\prime}-r}{B}+1)(e^{\epsilon}-1)+d^{\prime}}\cdot\frac{1}{rd^{\prime}(d^{\prime}-1)}$.
Similarly, the support matrix of the second configuration, $S^{(2)}$, is constructed from rows where $z>r$, resulting in a constant support size of $\lfloor\frac{d^{\prime}}{B}\rfloor$. Its base probability is computed as $p_{2}=\frac{d^{\prime}}{\frac{d^{\prime}-r}{B}(e^{\epsilon}-1)+d^{\prime}}\cdot\frac{1}{(B-r)d^{\prime}(d^{\prime}-1)}$. It can be verified that $p_{\alpha}p_{1}=p_{0}$ and $(1-p_{\alpha})p_{2}=p_{0}$, demonstrating that the sampling of these two configurations generates the same support matrix $S$.
These split configurations possess constant support sizes and extremal configurations. We now prove that their support matrix and base probability satisfy the requirements of a symmetric configuration. Taking $S^{(1)}$ as an example, this requires $p_{1}\sum_{o}S^{(1)}_{o,i}S^{(1)}_{o,j}$ to be constant for all $i\neq j$. For an arbitrary pair $(i,j)$ where $i\neq j$, the term $\sum_{o}S^{(1)}_{o,i}S^{(1)}_{o,j}$ represents the number of hash functions mapping $(i,j)$ into the same bucket of size $\lceil\frac{d^{\prime}}{B}\rceil$. Elements within such a bucket can form $(\lceil\frac{d^{\prime}}{B}\rceil+1)(\lceil\frac{d^{\prime}}{B}\rceil)$ distinct pairs, and there are $r$ such buckets. Mapping to the same bucket is equivalent to mapping to these distinct pairs. By Lemma D.1, there are exactly $r(\lceil\frac{d^{\prime}}{B}\rceil+1)(\lceil\frac{d^{\prime}}{B}\rceil)$ hash functions mapping $(i,j)$ to these pairs. This holds for every $(i,j)$, ensuring the sum is constant.
@@
Furthermore, extending this counting method to $B$ buckets and ignoring the remainder $r$, one can derive $P(h(x_{1})=h(x_{2}))\approx\frac{d^{\prime}-B}{B^{2}(d^{\prime}-1)}$, verifying the reduction in collision probability discussed in Section 5.2.
Finally, if a process generates variable $V_{1}$ with probability $p$ and variable $V_{2}$ with probability 1p1-p, and $E[V_{1}]=E[V_{2}]$, the variance of the process is $p\text{Var}(V_{1})+(1-p)\text{Var}(V_{2})$. This proves Corollary 5.1.
@@
where $\text{Var}(\hat{f}(x)\mid d^{\prime})$ denotes the variance calculated with $d^{\prime}$ in place of $d$. Assuming an optimal estimator (with the $k$ term addressed in the subsequent step), we set $k=\frac{d^{\prime}}{e^{\epsilon}+1}$ to optimize $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. Given Eq. (36), we obtain:
@@
Next, we evaluate the deviation of $k$ from the optimal value $k^{*}$ for a given dictionary size $d^{\prime}$. If $B=\frac{d^{\prime}}{k^{*}}$ is an integer, OCMS is equivalent to sampling between the configurations with support size of $\lceil k^{*}\rceil$ and $\lfloor k^{*}\rfloor$. If $B=\frac{d^{\prime}}{k^{*}}$ is not an integer, then $B$ is either $\lceil\frac{d^{\prime}}{k^{*}}\rceil$ or $\lfloor\frac{d^{\prime}}{k^{*}}\rfloor$, which correspond to sampling between configurations with support size of either ($k^{*}-1$ and $k^{*}$) or ($k^{*}+1$ and $k^{*}$). In all such scenarios, the support size of the sampled configurations deviates from $k^{*}$ by at most one. Analogous to Eq. (37), we define the relative change in variance as: